# Manifold-aware Representation Learning for Degradation-agnostic Image Restoration

### v1

@@ 4.1 Mixed Degradation Adaptation Block for AnyIR
Compared to standard attention-only models where $F_{\text{out}}^{\text{att}}$ is essentially a linear combination modulated by Softmax weights, the fused output in Alg. 1 introduces richer non-linear interactions. This enhances the model’s ability to fit complex degradation mappings, making it more suitable for real-world scenarios with mixed or ambiguous degradations. Subsequently, layer normalization, a feed-forward network (FFN), and a residual connection are applied: $F_{\text{out}}=\operatorname{FFN}(\operatorname{Norm}(F_{\text{out}}^{\text{fuse}}))+F_{\text{out}}^{\text{fuse}}$. This sequence stabilizes feature distributions and further boosts expressiveness.
@@
4.2 Shallow-Latent Contrastive Learning via SPD Manifold Alignment
Shallow-Latent Feature Pairs are Naturally Contrastive Pairs. Features extracted at different depths exhibit fundamentally different statistical properties. As shown in Fig. 4, shallow-stage features (e.g., Scale1) present sparse and decorrelated channel distributions, while deeper layers (e.g., Scale4) become increasingly redundant and concentrated. This trend is quantitatively supported by the effective rank ratio across scales, which increases from only 4.2% ($\nicefrac{{1}}{{24}}$ at 1st Scale) to 16.1% ($\nicefrac{{31}}{{192}}$ at 4th Scale). However, by compressing the deep features through a lightweight MLP, we obtain a latent representation with a significantly higher rank ratio of 36.5% ($\nicefrac{{35}}{{96}}$), indicating a more decorrelated and expressive embedding. This structural disparity between sparse, localized shallow features and compressed, semantic latent ones naturally defines a contrastive pairing without requiring additional augmentation. We leverage this depth-asymmetric contrast to impose consistency across stages, enabling better semantic alignment and stronger representational generalization under complex degradation conditions. Note that this case study is conducted under noise degradation; however, similar trends are consistently observed for other degradations as well. Please refer to our Supplementary Materials (i.e., Supp. Mat.) for more details.
SPD Manifold Space Contrastive Learning Leads to More Discriminative Representations. To enhance representation consistency across depth, we introduce a contrastive objective defined over SPD (Symmetric Positive Definite) manifold features. Given shallow features $F_{\text{shallow}}\in\mathbb{R}^{C_{s}\times H\times W}$ and latent features $F_{\text{latent}}\in\mathbb{R}^{C_{l}\times H^{\prime}\times W^{\prime}}$, we first reduce their channel dimensions via $1\times 1$ convolutions. The resulting tensors are reshaped into feature matrices $X_{s},X_{l}\in\mathbb{R}^{C\times N}$ with $N=H\times W$, and their second-order statistics are computed as:
@@ 4.2 Shallow-Latent Contrastive Learning via SPD Manifold Alignment
where $\mu$ is the mean across spatial dimensions, and $\epsilon\mathbf{I}$ ensures numerical stability and positive definiteness. The SPD matrices $\mathbf{C}_{s},\mathbf{C}_{l}\in\mathbb{R}^{C\times C}$ are vectorized and projected to a contrastive embedding space via shallow MLPs:
@@
where $\operatorname{sim}(\cdot,\cdot)$ denotes cosine similarity, and $\tau$ is a temperature parameter. Unlike conventional Euclidean contrastive learning which treats feature vectors as flat points our SPD-based method preserves second-order channel dependencies, enabling richer structural supervision. This manifold-aware regularization aligns local and semantic features across depth, enhances discriminability, and introduces no additional inference cost.
@@ 5 Experiments
We conduct experiments adhering to the protocols of prior general image restoration works potlapalli2023promptir ; zhang2023ingredient under four settings: (i) All-in-One (3Degradations), (ii) All-in-One (5Degradations), (iii) Mixed Degradation Setting, (iv) Adverse Weather Removal Setting, and (v) Zero-Shot Setting. More experimental details and the dataset introduction are provided in our Supp. Mat.
@@
3 Degradations. We evaluate our all-in-one restorer, MIRAGE , against other specialized methods listed in Tab. 1, all trained on three degradations: dehazing, deraining, and denoising. MIRAGE consistently outperforms all the comparison methods, even for those with the assistance of language, multi-task, or prompts. Notably, even our 6M tiny model outperforms the baseline method PromptIR potlapalli2023promptir by 0.71dB on average. And the 10M small model achieves the best performance across all the metrics, while maintaining 60% fewer parameters compared to MoCE-IR zamfir2025moce .
@@
BRDNet tian2000brdnet Neural Networks’22 - 23.23 .895 27.42 .895 32.26 .898 29.76 .836 26.34 .693 27.80 .843
LPNet gao2019dynamic CVPR’19 - 20.84 .828 24.88 .784 26.47 .778 24.77 .748 21.26 .552 23.64 .738
FDGAN dong2020fdgan AAAI’20 - 24.71 .929 29.89 .933 30.25 .910 28.81 .868 26.43 .776 28.02 .883
DL fan2019dl TPAMI’19 2M 26.92 .931 32.62 .931 33.05 .914 30.41 .861 26.90 .740 29.98 .876
MPRNet zamir2021multi CVPR’21 16M 25.28 .955 33.57 .954 33.54 .927 30.89 .880 27.56 .779 30.17 .899
AirNet li2022all CVPR’22 9M 27.94 .962 34.90 .967 33.92 .933 31.26 .888 28.00 .797 31.20 .910
NDR yao2024neural TIP’24 28M 25.01 .860 28.62 .848 28.72 .826 27.88 .798 26.18 .720 25.01 .810
PromptIR potlapalli2023promptir NeurIPS’23 36M 30.58 .974 36.37 .972 33.98 .933 31.31 .888 28.06 .799 32.06 .913
MoCE-IR-S zamfir2025moce CVPR’25 11M 30.98 .979 38.22 .983 34.08 .933 31.42 .888 28.16 .798 32.57 .916
AdaIR cui2025adair ICLR’25 29M 31.06 .980 38.64 .983 34.12 .935 31.45 .892 28.19 .802 32.69 .918
MoCE-IR zamfir2025moce CVPR’25 25M 31.34 .979 38.57 .984 34.11 .932 31.45 .888 28.18 .800 32.73 .917
MIRAGE -T (Ours) 2025 6M 31.81 .982 38.44 .983 34.05 .935 31.40 .892 28.14 .802 32.77 .919
MIRAGE -S (Ours) 2025 10M 31.86 .981 38.94 .985 34.12 .935 31.46 .891 28.19 .803 32.91 .919
@@
DA-CLIP luo2023controlling ICLR’24 125M 29.46 .963 36.28 .968 30.02 .821 24.86 .585 22.29 .476 - -
ArtPromptIR wu2024harmony ACM MM’24 36M 30.83 .979 37.94 .982 34.06 .934 31.42 .891 28.14 .801 32.49 .917
InstructIR-3D conde2024high ECCV’24 16M 30.22 .959 37.98 .978 34.15 .933 31.52 .890 28.30 .804 32.43 .913
UniProcessor duan2025uniprocessor ECCV’24 1002M 31.66 .979 38.17 .982 34.08 .935 31.42 .891 28.17 .803 32.70 .918
VLU-Net zeng2025vision CVPR’25 35M 30.71 .980 38.93 .984 34.13 .935 31.48 .892 28.23 .804 32.70 .919
@@
NAFNet chen2022simple ECCV’22 17M 25.23 .939 35.56 .967 31.02 .883 26.53 .808 20.49 .809 27.76 .881
DGUNet mou2022deep CVPR’22 17M 24.78 .940 36.62 .971 31.10 .883 27.25 .837 21.87 .823 28.32 .891
SwinIR liang2021swinir ICCVW’21 1M 21.50 .891 30.78 .923 30.59 .868 24.52 .773 17.81 .723 25.04 .835
Restormer zamir2022restormer CVPR’22 26M 24.09 .927 34.81 .962 31.49 .884 27.22 .829 20.41 .806 27.60 .881
MambaIR guo2024mambair ECCV’24 27M 25.81 .944 36.55 .971 31.41 .884 28.61 .875 22.49 .832 28.97 .901
DL fan2019dl TPAMI’19 2M 20.54 .826 21.96 .762 23.09 .745 19.86 .672 19.83 .712 21.05 .743
Transweather valanarasu2022transweather CVPR’22 38M 21.32 .885 29.43 .905 29.00 .841 25.12 .757 21.21 .792 25.22 .836
TAPE liu2022tape ECCV’22 1M 22.16 .861 29.67 .904 30.18 .855 24.47 .763 18.97 .621 25.09 .801
AirNet li2022all CVPR’22 9M 21.04 .884 32.98 .951 30.91 .882 24.35 .781 18.18 .735 25.49 .847
IDR zhang2023ingredient CVPR’23 15M 25.24 .943 35.63 .965 31.60 .887 27.87 .846 21.34 .826 28.34 .893
PromptIR potlapalli2023promptir NeurIPS’23 36M 26.54 .949 36.37 .970 31.47 .886 28.71 .881 22.68 .832 29.15 .904
MoCE-IR-S zamfir2025moce CVPR’25 11M 31.33 .978 37.21 .978 31.25 .884 28.90 .877 21.68 .851 30.08 .913
AdaIR cui2025adair ICLR’25 29 30.53 .978 38.02 .981 31.35 .889 28.12 .858 23.00 .845 30.20 .910
MoCE-IR zamfir2025moce CVPR’25 25M 30.48 .974 38.04 .982 31.34 .887 30.05 .899 23.00 .852 30.58 .919
MIRAGE -T (Ours) 2025 6M 31.35 .979 38.24 .983 31.35 .891 27.98 .850 23.11 .854 30.41 .912
MIRAGE -S (Ours) 2025 10M 31.45 .980 38.92 .985 31.41 .892 28.10 .858 23.59 .858 30.68 .914
@@
InstructIR-5D conde2024high ECCV’24 16M 36.84 .973 27.10 .956 31.40 .887 29.40 .886 23.00 .836 29.55 .908
ArtPromptIR wu2024harmony ACM MM’24 36M 29.93 .908 22.09 .891 29.43 .843 25.61 .776 21.99 .811 25.81 .846
VLU-Net zeng2025vision CVPR’25 35M 30.84 .980 38.54 .982 31.43 .891 27.46 .840 22.29 .833 30.11 .905
5 Degradations. Extending the three degradation tasks to include deblurring and low-light enhancement li2022all ; zhang2023ingredient , we validate our method’s comprehensive performance in an All-in-One setting. As shown in Tab. 2, MIRAGE -S effectively leverages degradation-specific features, surpassing PromptIR potlapalli2023promptir , MoCE-IR-S zamfir2025moce , AdaIR cui2025adair , and VLU-Net zeng2025vision by an average of 1.53dB, 0.6dB, 0.48dB, and 0.57dB with lower parameters. Notably, our tiny model (6M) also achieves competitive results (a second-best average PSNR) compared to the MoCE-IR (25M) and outperforms all other comparison methods, even those with the assistance of other modalities, multi-task learning, or pretraining.
Composited Degradation Setting.
@@
AirNet li2022all 9M 24.83 .778 24.21 .951 26.55 .891 26.79 .919 23.23 .779 22.82 .710 23.29 .723 22.21 .868 23.29 .901 21.80 .708 22.24 .725 23.75 .814
PromptIR potlapalli2023promptir 36M 26.32 .805 26.10 .969 31.56 .946 31.53 .960 24.49 .789 25.05 .771 24.51 .761 24.54 .924 23.70 .925 23.74 .752 23.33 .747 25.90 .850
WGWSNet ZhuWFYGDQH23 26M 24.39 .774 27.90 .982 33.15 .964 34.43 .973 24.27 .800 25.06 .772 24.60 .765 27.23 .955 27.65 .960 23.90 .772 23.97 .771 26.96 .863
WeatherDiff weather_diffusion 83M 23.58 .763 21.99 .904 24.85 .885 24.80 .888 21.83 .756 22.69 .730 22.12 .707 21.25 .868 21.99 .868 21.23 .716 21.04 .698 22.49 .799
OneRestore guo2024onerestore 6M 26.48 .826 32.52 .990 33.40 .964 34.31 .973 25.79 .822 25.58 .799 25.19 .789 29.99 .957 30.21 .964 24.78 .788 24.90 .791 28.47 .878
MoCE-IR zamfir2025moce 11M 27.26 .824 32.66 .990 34.31 .970 35.91 .980 26.24 .817 26.25 .800 26.04 .793 29.93 .964 30.19 .970 25.41 .789 25.39 .790 29.05 .881
@@
To simulate more realistic restoration scenarios, we extend the setting from OneRestore guo2024onerestore by including not only diverse single degradations—rain, haze, snow, low illumination—but also composite degradations where multiple types are combined within the same image. This results in a total of eleven distinct restoration settings. As shown in Tab. 3, our Tiny (6M) and Small (10M) models outperform OneRestore guo2024onerestore (6M) by 0.39 dB and 0.86 dB on average, respectively. Compared to the recent state-of-the-art MoCE-IR zamfir2025moce (11M), our Small model still achieves 0.28 dB higher performance while being more compact (10M vs. 11M). These results demonstrate the effectiveness of our method, especially in handling complex, mixed degradations.
@@
Method Venue Snow100K-S Snow100K_dataset Snow100K-L Snow100K_dataset Outdoor-Rain test1_dataset_allweather RainDrop raindrop Average
@@
All-in-One as2020 CVPR’20 28.33 .882 24.71 .898 31.12 .927 28.05 .902
TransWeather Transweather CVPR’22 32.51 .934 29.31 .888 28.83 .900 30.17 .916 30.20 .909
Chen et al. ChenHTYDK22 CVPR’22 34.42 .947 30.22 .907 29.27 .915 31.81 .931 31.43 .925
WGWSNet ZhuWFYGDQH23 CVPR’23 34.31 .946 30.16 .901 29.32 .921 32.38 .938 31.54 .926
WeatherDiff64 weather_diffusion TPAMI’23 35.83 .957 30.09 .904 29.64 .931 30.71 .931 31.57 .931
WeatherDiff128 weather_diffusion TPAMI‘23 35.02 .952 29.58 .894 29.72 .922 29.66 .923 31.00 .923
AWRCP AWRCP_YeCBSXJYCL23 ICCV’23 36.92 .965 31.92 .934 31.39 .933 31.93 .931 33.04 .941
GridFormer Gridformer IJCV’24 37.46 .964 31.71 .923 31.87 .933 32.39 .936 33.36 .939
MPerceiver AiHZW024 CVPR’24 36.23 .957 31.02 .916 31.25 .925 33.21 .929 32.93 .932
DTPM DTPM_0001CCXQL024 CVPR’24 37.01 .966 30.92 .917 30.99 .934 32.72 .944 32.91 .940
Histoformer Histoformer_SunRGWC24 ECCV’24 37.41 .966 32.16 .926 32.08 .939 33.06 .944 33.68 .944
MIRAGE -S (Ours) 2025 37.97 .973 32.33 .929 32.82 .949 32.78 .945 33.98 .949
Adverse Weather Removal Setting. Following valanarasu2022transweather ; zhu2023learning , We test our MIRAGE on three challenging deweathering tasks: snow removal, rain streak and fog removal, and raindrop removal. Tab. 4 shows the comparison of our MIRAGE and other state-of-the-art methods. MIRAGE consistently outperforms existing methods across almost all datasets except the PSNR performance for RainDrop. The significant performance gains over multiple weather degradations demonstrate the effectiveness of MIRAGE in handling diverse weather-related degradations. Especially, 0.3 dB improvement on PSNR over Histoformer Histoformer_SunRGWC24 and 1.05 dB improvements over MPerceiver AiHZW024 .
@@
Method PSNR (dB, $\uparrow$) SSIM ($\uparrow$)
SwinIR liang2021swinir 15.31 .740
NAFNet chu2022nafssr 15.42 .744
Restormer zamir2022restormer 15.46 .745
AirNet li2022all 15.46 .745
IDR zhang2023ingredient 15.58 .762
PromptIR potlapalli2023promptir 15.48 .748
MoCE-IR zamfir2025moce 15.91 .765
@@
Method PSNR (dB, $\uparrow$) Memory ($\downarrow$) Params. ($\downarrow$) FLOPs ($\downarrow$)
AirNet li2022all 31.20 4829M 8.93M 238G
PromptIR potlapalli2023promptir 32.06 9830M 35.59M 132G
IDR zhang2023ingredient - 4905M 15.34M 98G
AdaIR cui2025adair - 9740M 28.79M 124G
MoCE-IR-S zamfir2025moce 32.51 4263M 11.48M 37G
MoCE-IR zamfir2025moce 32.73 6654M 25.35M 75G

### v2

@@ 4.1 Mixed Degradation Adaptation Block for Degradation-Agnostic IR
Compared to the attention-only models, the fused output in Alg. 1 introduces richer interactions. This enhances the model’s ability to fit complex degradation mappings, making it more suitable for mixed or ambiguous degradations. Subsequently, layer normalization, a feed-forward network (FFN), and a residual connection are applied: $F_{\text{out}}=\operatorname{FFN}(\operatorname{Norm}(F_{\text{out}}^{\text{fuse}}))+F_{\text{out}}^{\text{fuse}}$. This sequence stabilizes feature distributions and further boosts expressiveness.
4.2 Shallow-Latent Contrastive Learning via SPD Manifold Alignment
The unified IR model requires a single backbone to process degradations that depend on fundamentally different representational levels. Shallow layers primarily encode degradation-specific, fine-grained structures, whereas deeper layers become more semantic and statistically stable. This inherent depth asymmetry introduces representation drift when multiple degradations share the same feature space, motivating a mechanism that explicitly enforces cross-stage consistency. We therefore treat shallow and latent features as complementary views of the underlying signal and align them to stabilize the shared representation space, thereby improving generalization across heterogeneous degradations.
Shallow-Latent Feature Pairs are Naturally Contrastive Pairs. Features extracted at different depths exhibit fundamentally different statistical properties. As shown in Fig. 4, shallow-stage features (e.g., Scale1) present sparse and decorrelated channel distributions, while deeper layers (e.g., Scale4) become increasingly redundant and concentrated. This trend is quantitatively supported by the effective rank ratio across scales, which increases from only 4.2% ($\nicefrac{{1}}{{24}}$ at 1st Scale) to 16.1% ($\nicefrac{{31}}{{192}}$ at 4th Scale). However, by compressing the deep features through a lightweight MLP, we obtain a latent representation with a notably higher rank ratio of 36.5% ($\nicefrac{{35}}{{96}}$), indicating a more decorrelated and expressive embedding. This structural disparity between sparse, localized shallow features and compressed, semantic latent ones naturally defines a contrastive pairing without requiring augmentation. We leverage this depth-asymmetric contrast to impose consistency across stages, enabling better semantic alignment and stronger representational generalization under complex degradation conditions. Note that this study is conducted under noise degradation; however, similar trends are consistently observed for other degradations as well. See the appendix for more details.
@@ 4.2 Shallow-Latent Contrastive Learning via SPD Manifold Alignment
SPD Manifold Space Contrastive Learning Leads to More Discriminative Representations. To enhance representation consistency across depth, we introduce a contrastive objective defined over SPD (Symmetric Positive Definite) manifold features. We note that the goal here is not to perform full Riemannian optimization along SPD geodesics. Instead, we adopt a lightweight formulation that retains the key second‐order structure of covariance matrices while keeping training stable and efficient. Strict geodesic contrastive learning typically requires repeated log/exp mappings and matrix decompositions, which incur considerable overhead in large low-level vision models. Our approach strikes a practical balance by preserving essential SPD structure before projection. Specifically, in our work, given shallow features $F_{\text{shallow}}\in\mathbb{R}^{C_{s}\times H\times W}$ and latent features $F_{\text{latent}}\in\mathbb{R}^{C_{l}\times H^{\prime}\times W^{\prime}}$, we first reduce their channel dimensions via $1\times 1$ convolutions. The resulting tensors are reshaped into feature matrices $X_{s},X_{l}\in\mathbb{R}^{C\times N}$ with $N=H\times W$, and their second-order statistics are computed as:
@@
where $\mu$ is the mean across spatial dimensions, and $\epsilon\mathbf{I}$ ensures numerical stability and positive definiteness. The SPD matrices $\mathbf{C}_{s},\mathbf{C}_{l}\in\mathbb{R}^{C\times C}$ are vectorized and projected to a contrastive embedding space via shallow 1-layer MLPs:
@@
where $\operatorname{sim}(\cdot,\cdot)$ denotes cosine similarity and $\tau$ a temperature parameter. Unlike Euclidean contrastive learning, which views features as flat vectors, our SPD-based method preserves second-order channel dependencies, providing richer structural supervision. This regularization aligns local and semantic features across depth, enhances discriminability, and introduces no additional inference cost.
@@ 5 Experiments
We conduct experiments adhering to the protocols of prior general image restoration works (Potlapalli et al., 2024; Zhang et al., 2023) under 5 settings: (a) 3 Degradations), (b) 5 Degradations), (c) Mixed Degradation, (d) Adverse Weather Removal, and (e) Zero-Shot. The implementation and experimental details, and the dataset description are provided in the appendix. Our code, checkpoints, and visual results are available via: https://github.com/Amazingren/MIRAGE.
@@
3 Degradations. We evaluate our method against others listed in Tab. 1, all trained on three degradations: dehazing, deraining, and denoising. MIRAGE consistently outperforms all the comparison methods, even for those with the assistance of language, multi-task, or prompts. Notably, even our 6M tiny model outperforms our baseline PromptIR by 0.71dB on average. Our 10M small model achieves the best performance across all the metrics, with 60% fewer parameters compared MoCE-IR. Compared to DA-RCOT (Tang et al., 2025b), which performs contrastive learning over residual feature space, MIRAGE achieves consistently better restoration quality while using substantially fewer parameters (10M vs. 50M). This highlights the efficiency and effectiveness of our SPD-based cross-layer alignment despite its more compact design.
@@
BRDNet (Tian et al., 2020) NN’20 - 23.23 .895 27.42 .895 32.26 .898 29.76 .836 26.34 .693 27.80 .843
LPNet (Gao et al., 2019) CVPR’19 - 20.84 .828 24.88 .784 26.47 .778 24.77 .748 21.26 .552 23.64 .738
FDGAN (Dong et al., 2020) AAAI’20 - 24.71 .929 29.89 .933 30.25 .910 28.81 .868 26.43 .776 28.02 .883
DL (Fan et al., 2019) TPAMI’19 2M 26.92 .931 32.62 .931 33.05 .914 30.41 .861 26.90 .740 29.98 .876
MPRNet (Zamir et al., 2021) CVPR’21 16M 25.28 .955 33.57 .954 33.54 .927 30.89 .880 27.56 .779 30.17 .899
AirNet (Li et al., 2022) CVPR’22 9M 27.94 .962 34.90 .967 33.92 .933 31.26 .888 28.00 .797 31.20 .910
NDR (Yao et al., 2024) TIP’24 28M 25.01 .860 28.62 .848 28.72 .826 27.88 .798 26.18 .720 25.01 .810
PromptIR (Potlapalli et al., 2024) NeurIPS’23 36M 30.58 .974 36.37 .972 33.98 .933 31.31 .888 28.06 .799 32.06 .913
MoCE-IR-S (Zamfir et al., 2025) CVPR’25 11M 30.98 .979 38.22 .983 34.08 .933 31.42 .888 28.16 .798 32.57 .916
AdaIR (Cui et al., 2025) ICLR’25 29M 31.06 .980 38.64 .983 34.12 .935 31.45 .892 28.19 .802 32.69 .918
MoCE-IR (Zamfir et al., 2025) CVPR’25 25M 31.34 .979 38.57 .984 34.11 .932 31.45 .888 28.18 .800 32.73 .917
DA-RCOT (Tang et al., 2025b) TPAMI’25 50M 31.26 .977 38.36 .983 33.98 .934 31.33 .890 28.10 .801 32.60 .917
MIRAGE -T (Ours) ICLR’26 6M 31.81 .982 38.44 .983 34.05 .935 31.40 .892 28.14 .802 32.77 .919
MIRAGE -S (Ours) ICLR’26 10M 31.86 .981 38.94 .985 34.12 .935 31.46 .891 28.19 .803 32.91 .919
@@
DA-CLIP (Luo et al., 2024) ICLR’24 125M 29.46 .963 36.28 .968 30.02 .821 24.86 .585 22.29 .476 - -
ArtPromptIR (Wu et al., 2024) ACM MM’24 36M 30.83 .979 37.94 .982 34.06 .934 31.42 .891 28.14 .801 32.49 .917
InstructIR-3D (Conde et al., 2024) ECCV’24 16M 30.22 .959 37.98 .978 34.15 .933 31.52 .890 28.30 .804 32.43 .913
UniProcessor (Duan et al., 2025) ECCV’24 1002M 31.66 .979 38.17 .982 34.08 .935 31.42 .891 28.17 .803 32.70 .918
VLU-Net (Zeng et al., 2025) CVPR’25 35M 30.71 .980 38.93 .984 34.13 .935 31.48 .892 28.23 .804 32.70 .919
RamIR (Tang et al., 2025a) Applied’25 21.7M 31.29 .977 38.16 .981 34.04 .931 31.61 .891 28.19 .801 32.65 .916
@@
NAFNet (Chen et al., 2022a) ECCV’22 17M 25.23 .939 35.56 .967 31.02 .883 26.53 .808 20.49 .809 27.76 .881
DGUNet (Mou et al., 2022) CVPR’22 17M 24.78 .940 36.62 .971 31.10 .883 27.25 .837 21.87 .823 28.32 .891
SwinIR (Liang et al., 2021) ICCVW’21 1M 21.50 .891 30.78 .923 30.59 .868 24.52 .773 17.81 .723 25.04 .835
Restormer (Zamir et al., 2022) CVPR’22 26M 24.09 .927 34.81 .962 31.49 .884 27.22 .829 20.41 .806 27.60 .881
MambaIR (Guo et al., 2024a) ECCV’24 27M 25.81 .944 36.55 .971 31.41 .884 28.61 .875 22.49 .832 28.97 .901
DL (Fan et al., 2019) TPAMI’19 2M 20.54 .826 21.96 .762 23.09 .745 19.86 .672 19.83 .712 21.05 .743
Transweather CVPR’22 38M 21.32 .885 29.43 .905 29.00 .841 25.12 .757 21.21 .792 25.22 .836
TAPE (Liu et al., 2022) ECCV’22 1M 22.16 .861 29.67 .904 30.18 .855 24.47 .763 18.97 .621 25.09 .801
AirNet (Li et al., 2022) CVPR’22 9M 21.04 .884 32.98 .951 30.91 .882 24.35 .781 18.18 .735 25.49 .847
IDR (Zhang et al., 2023) CVPR’23 15M 25.24 .943 35.63 .965 31.60 .887 27.87 .846 21.34 .826 28.34 .893
PromptIR (Potlapalli et al., 2024) NeurIPS’23 36M 30.41 .972 36.17 .970 31.20 .885 27.93 .851 22.89 .829 29.72 .901
MoCE-IR-S (Zamfir et al., 2025) CVPR’25 11M 31.33 .978 37.21 .978 31.25 .884 28.90 .877 21.68 .851 30.08 .913
AdaIR (Cui et al., 2025) ICLR’25 29 30.53 .978 38.02 .981 31.35 .889 28.12 .858 23.00 .845 30.20 .910
MoCE-IR (Zamfir et al., 2025) CVPR’25 25M 30.48 .974 38.04 .982 31.34 .887 30.05 .899 23.00 .852 30.58 .919
DA-RCOT (Tang et al., 2025b) TPAMI’25 50M 30.96 .975 37.87 .980 31.23 .888 28.68 .872 23.25 .836 30.40 .911
MIRAGE -T (Ours) ICLR’26 6M 31.35 .979 38.24 .983 31.35 .891 27.98 .850 23.11 .854 30.41 .912
MIRAGE -S (Ours) ICLR’26 10M 31.45 .980 38.92 .985 31.41 .892 28.10 .858 23.59 .858 30.68 .914
@@
InstructIR-5D (Conde et al., 2024) ECCV’24 16M 36.84 .973 27.10 .956 31.40 .887 29.40 .886 23.00 .836 29.55 .908
ArtPromptIR (Wu et al., 2024) ACM MM’24 36M 29.93 .908 22.09 .891 29.43 .843 25.61 .776 21.99 .811 25.81 .846
VLU-Net (Zeng et al., 2025) CVPR’25 35M 30.84 .980 38.54 .982 31.43 .891 27.46 .840 22.29 .833 30.11 .905
RamIR (Tang et al., 2025a) Applied’25 21.7M 31.09 .978 37.56 .979 31.44 .886 28.82 .878 22.02 .828 30.18 .910
5 Degradations. Extending the 3 tasks with deblurring and low-light enhancement (Li et al., 2022; Zhang et al., 2023), we evaluate our MIRAGE ’s performance in a more challenging 5-degradation setting. Tab. 2 shows that MIRAGE -S surpasses PromptIR (Potlapalli et al., 2024), MoCE-IR-S (Zamfir et al., 2025), AdaIR (Cui et al., 2025), and VLU-Net (Zeng et al., 2025) by 1.53dB, 0.6dB, 0.48dB, and 0.57dB on average, with fewer parameters. Our tiny model (6M) also achieves a second-best average PSNR against MoCE-IR (25M) and surpasses all other methods, including those aided by additional modalities, multi-task learning, or pretraining.
Mixed Degradations.
@@
AirNet 9M 24.83 .778 24.21 .951 26.55 .891 26.79 .919 23.23 .779 22.82 .710 23.29 .723 22.21 .868 23.29 .901 21.80 .708 22.24 .725 23.75 .814
PromptIR 36M 26.32 .805 26.10 .969 31.56 .946 31.53 .960 24.49 .789 25.05 .771 24.51 .761 24.54 .924 23.70 .925 23.74 .752 23.33 .747 25.90 .850
WGWSNet 26M 24.39 .774 27.90 .982 33.15 .964 34.43 .973 24.27 .800 25.06 .772 24.60 .765 27.23 .955 27.65 .960 23.90 .772 23.97 .771 26.96 .863
WeatherDiff 83M 23.58 .763 21.99 .904 24.85 .885 24.80 .888 21.83 .756 22.69 .730 22.12 .707 21.25 .868 21.99 .868 21.23 .716 21.04 .698 22.49 .799
OneRestore 6M 26.48 .826 32.52 .990 33.40 .964 34.31 .973 25.79 .822 25.58 .799 25.19 .789 29.99 .957 30.21 .964 24.78 .788 24.90 .791 28.47 .878
MoCE-IR 11M 27.26 .824 32.66 .990 34.31 .970 35.91 .980 26.24 .817 26.25 .800 26.04 .793 29.93 .964 30.19 .970 25.41 .789 25.39 .790 29.05 .881
@@
To better approximate real-world conditions, we extend OneRestore (Guo et al., 2024b) to cover both diverse single degradations (rain, haze, snow, low light) and composite cases with multiple degradations per image, yielding eleven distinct restoration settings. As shown in Tab. 3, MIRAGE consistently outperforms leading approaches including AirNet (Li et al., 2022), PromptIR (Potlapalli et al., 2024), WGWSNet (Zhu et al., 2023), WeatherDiff (Özdenizci and Legenstein, 2023), OneRestore (Guo et al., 2024b), and MoCE-IR (Zamfir et al., 2025). Specifically, our Tiny (6M) and Small (10M) models outperform OneRestore (Guo et al., 2024b) (6M) by 0.39 dB and 0.86dB on average. Compared to the recent SOTA MoCE-IR (Zamfir et al., 2025) (11M), our Small model achieves 0.28dB higher performance with fewer parameters (10M vs. 11M). These results highlight the effectiveness of our method, particularly for complex mixed degradations.
@@
Method Venue Snow100K-S Snow100K-L Outdoor-Rain RainDrop Average
@@
All-in-One (Li et al., 2020) CVPR’20 28.33 .882 24.71 .898 31.12 .927 28.05 .902
TransWeather (Valanarasu et al., 2022) CVPR’22 32.51 .934 29.31 .888 28.83 .900 30.17 .916 30.20 .909
Chen et al. (Chen et al., 2022b) CVPR’22 34.42 .947 30.22 .907 29.27 .915 31.81 .931 31.43 .925
WGWSNet (Zhu et al., 2023) CVPR’23 34.31 .946 30.16 .901 29.32 .921 32.38 .938 31.54 .926
WeatherDiff64 (Özdenizci and Legenstein, 2023) TPAMI’23 35.83 .957 30.09 .904 29.64 .931 30.71 .931 31.57 .931
WeatherDiff128 (Özdenizci and Legenstein, 2023) TPAMI‘23 35.02 .952 29.58 .894 29.72 .922 29.66 .923 31.00 .923
AWRCP (Ye et al., 2023) ICCV’23 36.92 .965 31.92 .934 31.39 .933 31.93 .931 33.04 .941
GridFormer (Wang et al., 2024) IJCV’24 37.46 .964 31.71 .923 31.87 .933 32.39 .936 33.36 .939
MPerceiver (Ai et al., 2024) CVPR’24 36.23 .957 31.02 .916 31.25 .925 33.21 .929 32.93 .932
DTPM (Ye et al., 2024) CVPR’24 37.01 .966 30.92 .917 30.99 .934 32.72 .944 32.91 .940
Histoformer (Sun et al., 2024) ECCV’24 37.41 .966 32.16 .926 32.08 .939 33.06 .944 33.68 .944
MIRAGE -S (Ours) ICLR’26 37.97 .973 32.33 .929 32.82 .949 32.78 .945 33.98 .949
Adverse Weather Removal. Following (Valanarasu et al., 2022; Zhu et al., 2023), We test our MIRAGE on three challenging deweathering tasks: snow removal, rain streak and fog removal, and raindrop removal. Tab. 4 shows the comparison of our MIRAGE and other state-of-the-art methods. MIRAGE consistently outperforms existing methods across almost all datasets except PSNR for RainDrop. The performance gains over multiple weather degradations demonstrate the effectiveness of MIRAGE in handling diverse weather conditions. Especially, 0.30dB improvement on PSNR over Histoformer (Sun et al., 2024) and 1.05dB improvements over MPerceiver (Ai et al., 2024).
@@
Method PSNR ($\uparrow$) SSIM ($\uparrow$)
SwinIR (Liang et al., 2021) 15.31 .740
NAFNet (Chu et al., 2022) 15.42 .744
Restormer (Zamir et al., 2022) 15.46 .745
AirNet (Li et al., 2022) 15.46 .745
IDR (Zhang et al., 2023) 15.58 .762
PromptIR (Potlapalli et al., 2024) 15.48 .748
MoCE-IR (Zamfir et al., 2025) 15.91 .765
@@
Method PSNR ($\uparrow$) Memory ($\downarrow$) Params. ($\downarrow$) FLOPs ($\downarrow$)
AirNet (Li et al., 2022) 31.20 4829M 8.93M 238G
PromptIR (Potlapalli et al., 2024) 32.06 9830M 35.59M 132G
IDR (Zhang et al., 2023) - 4905M 15.34M 98G
AdaIR (Cui et al., 2025) - 9740M 28.79M 124G
MoCE-IR-S (Zamfir et al., 2025) 32.51 4263M 11.48M 37G
MoCE-IR (Zamfir et al., 2025) 32.73 6654M 25.35M 75G